OpenStudy (narissa):

which point represents a solution to the system x-y >( with line under that symbol) 7 x+2y<5

OpenStudy (narissa):

@Photon336

OpenStudy (narissa):

@ChantySquirrel1129**

OpenStudy (photon336):

@ganeshie8

OpenStudy (narissa):

@malcolmmcswain

OpenStudy (narissa):

@paki

OpenStudy (chantysquirrel1129**):

any answer choices? c:

OpenStudy (narissa):

yes 6,1 8,-1 6,2 6,-2

Nnesha (nnesha):

graph) or test each option by plugging in x and y values into both inequalities and see which one comes out as a true statement

OpenStudy (narissa):

im thinking its 6,-2

OpenStudy (narissa):

oops i meant 6,2

Nnesha (nnesha):

how do you know ?

OpenStudy (narissa):

7-5 is 2

Nnesha (nnesha):

\[\rm x-y \ge 7\] this the first inequality substitute (6,2) for x and y do you get true statement at the end ? :=))

OpenStudy (narissa):

ooo no

OpenStudy (narissa):

8,-1

Nnesha (nnesha):

how ?? post the work remember you should test by plugging into *both* inequality

OpenStudy (narissa):

k im going to try thanks

OpenStudy (narissa):

what does that symbol mean

OpenStudy (narissa):

same???

Nnesha (nnesha):

\[\ge \] means left side value MUST greater than or equal to right side in order to make a true statement \[x \ge 2\] x is ` greater than or equal` to 2 x can be 2 or greater than 2

OpenStudy (narissa):

its 6,-2

OpenStudy (narissa):

thank u

Nnesha (nnesha):

did you plugg into 2nd inequality bec the order pair should be true for both inequalities

OpenStudy (narissa):

im sooo confused

OpenStudy (narissa):

forget it its too hard i will wait i am in special ed

Nnesha (nnesha):

you have to inequalities \[\rm x-y \ge 7\] AND \[x+2y < 5\]you should substitute x and y into BOTH inequalities what do you get when (x,y)= did you replace x and y for 6,-2 into x+2y <5 ???

OpenStudy (narissa):

yaay but i just dont know

Nnesha (nnesha):

you don't know what ?

Nnesha (nnesha):

what did you get at left side ?? x+2y <5 substitute (X,y) for (6,-2) simplify left side what number would you get ?

OpenStudy (narissa):

6+2*-2right

Nnesha (nnesha):

yes what would be the final answer ? simplify that 6+2(-2) ?

OpenStudy (narissa):

2

Nnesha (nnesha):

so you get 2 at left side now is it true statement 2 <5 ???

OpenStudy (narissa):

ooooh wow i was thinking the other way the symbol oh gosh thanks for explaining

Nnesha (nnesha):

2 <5 means 2 is less than 5 is it true ??